# A Proof of Gross’ Conjecture on $2$-Automorphic $2$-Groups

### v1

@@ Abstract.
Building upon previous results, a classification is given of finite $p$-groups whose elements of order $p$ are all fused. This particularly confirms a conjecture of Gross proposed in 1976 on $2$-automorphic $2$-groups, which are $2$-groups with involutions forming a single fusion classes. As a consequence, two open problems regarding AT-groups and FIF-groups are solved.
@@ 1. Introduction
It is well-known that finite $p$-groups with a single subgroup of order $p$ are cyclic or a generalized quaternion group. The class of finite $p$-groups with a single fusion class of elements of order $p$ has been extensively studied. (Recall that a fusion class of a group $N$ is an orbit of ${\mathrm{Aut}}(N)$ on the elements.) The study was initiated by G. Higman 1963 when he classified finite non-abelian $2$-groups (with more than one involution) that has a cyclic group of automorphisms transitive on the involutions, called Suzuki $2$-groups.
For odd primes $p$, Shult [22] showed that homocyclic $p$-groups are the only $p$-groups of which elements of order $p$ form a single fusion class. Shaw [21] showed that a finite non-abelian $2$-group (with more than one involution) which has a solvable group of automorphisms transitive on the involutions is a Suzuki $2$-group. Higman’s classification associated with Shaw’s work forms Sections 5-7 of Huppert’s book [12, Chapter VIII].
@@
The proof of Theorem 1.1 for $2$-groups depends on certain representations of finite transitive linear groups over ${\mathbb{F}}_{2}$, stated below, which is independently interesting.
@@
Solving Gross’ conjecture is also an important part for the study of some other classes of groups. A group $N$ is called an AT-group if ${\mathrm{Aut}}(N)$ is transitive on the elements of the same order. Zhang [25] presents a description of finite AT-groups. However, the problem of determining AT-$2$-groups has been left open, as commented that “If Gross’ conjecture on $2$-automorphic $2$-groups is true, …… We do not know if Suzuki $2$-groups are AT-groups” before Lemma 2.2 of [25]. A group is called an FIF-group if any two elements of the same order are fused or inverse-fused. Praeger and the first-named author [16] give a description of finite FIF-groups, with the problem of classifying FIF-$2$-groups left open, see [16, Problem 1.5]. As a corollary of Theorem 1.1, these problems are solved.
@@
We remark that the automorphism groups of the groups $A_{2}(n,\theta)$, $B_{2}(n)$ and $P(\epsilon)$ are determined in Lemmas 4.2, 4.4 and 4.7.
@@ Acknowledgments
The authors acknowledge the support of NNSFC grant no. 11931005.
@@ 2. Transitive linear groups and their exterior squares
All fields and groups are finite in this section. For a field ${\mathrm{F}}$ and a group $G$, an ${\mathrm{F}}G$-reprensentation is a group homomorphism $\rho$ from $G$ to ${\mathrm{GL}}(V)$ for some ${\mathrm{F}}$-space $V$. Every ${\mathrm{F}}G$-reprensentation $\rho$ induces an ${\mathrm{F}}G$-module structure on $V$ by $(xg)\cdot v=x\cdot v^{\rho(g)}$ for $x\in{\mathrm{F}}$ and $g\in G$. If ${\mathrm{E}}$ is an extension field of ${\mathrm{F}}$, then $V^{\mathrm{E}}:=V\otimes{\mathrm{E}}$ is naturally an ${\mathrm{E}}G$-module. Some elementary properties of the operator $\cdot\otimes{\mathrm{E}}$ can be found in [12, Section VII.1], [1, Chapter 9] and [13, Chapter 9]. For an ${\mathrm{F}}G$-module $V$ with representation $\rho$, we use the following two notations:
@@
    [12, Chapter VII, 1.16(a)] $(V_{\mathrm{K}})^{\mathrm{F}}\cong\sum_{\sigma\in{\mathrm{Gal}}({\mathrm{F}}/{\mathrm{K}})}V^{\sigma}$;
@@
    [Deuring-Noether, see [12, Chapter VII, 1.22]] $V^{\mathrm{E}}\cong W^{\mathrm{E}}$ if and only if $V\cong W$.
@@
Let $V={\mathbb{F}}_{2}^{n}$ be a vector space, and let $G\leqslant{\mathrm{GL}}(V)\cong{\mathrm{GL}}_{n}(2)$ be transitive on $V\setminus\{0\}$. These groups $G$ are known by the classification of the finite $2$-transitive permutation groups, and we list them in the next theorem, see [4, Table 7.3].
@@
The following lemma can be directly deduced by Steinberg’s Twisted Tensor Product Theorem [23] (also see [14, Theorem 5.4.1]).
@@
Lemma 2.7 ([17, Theorem 2.2]).
Let $U={\mathbb{F}}_{q}^{m}$, and let $G={\mathrm{SL}}_{m}(q)$. Then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is an irreducible ${\mathbb{F}}_{q}G$-module which cannot be written over an subfield of ${\mathbb{F}}_{q}$.
@@
Results of exterior square of the natural modules of symplectic groups can be found in [5, 15, 19], and the result of the exceptional group of Lie type ${\mathrm{G}}_{2}(q)<{\mathrm{Sp}}_{6}(q)$ (even $q$) can be deduced from [20, Theorem 8.14].
@@
Suppose that $U={\mathbb{F}}_{q}^{2m}$ is a symplectic space with $q=2^{f}$, and $G={\mathrm{Sp}}_{2m}(q)$ or ${\mathrm{G}}_{2}(q)$($m=3$). Then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ has a maximal ${\mathbb{F}}_{q}G$-submodule $T$ of codimension $1$. Moerover, $T$ is indecomposable and one the following statements holds:
@@
Let $G\leqslant{\mathrm{GL}}(V)$ with $V={\mathbb{F}}_{2}^{n}$, and let $M=\Lambda^{2}_{{\mathbb{F}}_{2}}(V)/W$ for a submodule $W$. Assume that $G$ is transitive on $V\setminus\{0\}$. Then $G$ is listed in Theorem 2.4. If $G$ is one of the small groups listed in Theorem 2.4(ii), then computations in Magma [2] shows that $\dim M=1$. Thus, we only need to focus on the cases that $G^{(\infty)}\cong{\mathrm{SL}}_{m}(q)$ ($m\geqslant 2$ and $(m,f)\neq(2,1)$), ${\mathrm{Sp}}_{2m}(q)$ ($m\geqslant 2$ and $(m,q)\neq(2,2)$) or ${\mathrm{G}}_{2}(q)$ ($q\geqslant 4$) for $q=2^{f}$.
Let $K$ be the kernel of $G$ acting on $M$. Then $G/K$ is faithful and transitive on $M\setminus\{0\}$, and either $G^{(\infty)}\leqslant K$ or $K\leqslant{\mathbf{Z}}(G^{(\infty)})$. Suppose that $G^{(\infty)}\leqslant K$. Then $G/K$ is isomorphic to a quotient group of $G/G^{(\infty)}\lesssim{\Gamma\mathrm{L}}(1,2^{f})$. So
@@
We thus have that $K\leqslant{\mathbf{Z}}(G^{\infty})$. So $G/K$ is faithful and transitive on $M\setminus\{0\}$, and has a simple composition factor isomorphic to one of ${\mathrm{PSL}}_{m}(q)$ ($m\geqslant 2$ and $(m,f)\neq(2,1)$), ${\mathrm{PSp}}_{2m}(q)$ ($m\geqslant 2$ and $(m,q)\neq(2,2)$) and ${\mathrm{G}}_{2}(q)$ ($q\geqslant 4$) for even $q$. Note that none of these groups is isomorphic to ${\mathrm{A}}_{6}$, ${\mathrm{A}}_{7}$ or ${\mathrm{PSU}}_{3}(3)$, inspecting the candidates for $G/K$ in Theorem 2.4, we conclude that $(G/K)^{(\infty)}\cong G^{(\infty)}$. It follows that $K\cap G^{(\infty)}=1$ and $G^{(\infty)}$ is transitive on $M\setminus\{0\}$, and further $G^{(\infty)}$ is also transitive on $V\setminus\{0\}$.
@@
Lemma 2.6 deduces that each $B_{j}$ is an irreducible ${\mathbb{F}}_{2}G^{(\infty)}$-module. Suppose that $B_{j}$ is not a subspace of $W$ for some $j$. Then $M\cong B_{j}$ as ${\mathbb{F}}_{2}G^{(\infty)}$-modules. It follows that $\dim M=\dim B_{j}$. Note that $\dim B_{j}=n^{2}/f$ for $j\neq f/2$ and $\dim B_{f/2}=n^{2}/(2f)$ when $f$ is even. Hence, $M\cong B_{j}$ only when $j=f/2$, $f=n/2$ and $G^{(\infty)}\cong{\mathrm{SL}}_{2}(2^{f})$. In this case, the induced linear group of $G^{(\infty)}$ on $B_{f/2}$ is isomorphic to $\Omega^{-}_{4}(2^{f/2})$ (see [14, page 45]), which cannot be transitive on non-zero vectors of ${\mathbb{F}}_{2}^{2f}$. Thus, $B_{j}\leqslant W$ for each $j$, and then $M$ is ${\mathbb{F}}_{2}G^{(\infty)}$ isomorphic to a quotient of $A\cong(\Lambda^{2}_{{\mathbb{F}}_{q}}(U))_{{\mathbb{F}}_{2}}$. Suppose that $G^{(\infty)}\cong{\mathrm{Sp}}_{2m}(q)$ for $m\geqslant 2$ (or $G^{(\infty)}\cong G_{2}(q)$) with $q=2^{f}$. Let $T$ and $T_{0}$ be ${\mathbb{F}}_{q}G^{(\infty)}$-submodules of $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ give in Lemma 2.8. Assume that $m$ is even. Then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is indecomposable with $T$ a maximal ${\mathbb{F}}_{q}G^{(\infty)}$-submodule by Lemma 2.8. Note that $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)/T$ is a trivial ${\mathbb{F}}_{q}G^{(\infty)}$-module. This yields that $M$ is a trivial ${\mathbb{F}}_{2}G^{(\infty)}$-module, a contradiction. Assume that $m$ is odd. Then $T$ is irreducible and cannot be written over any subfield by Lemma 2.8. Hence, $T_{{\mathbb{F}}_{2}}$ is an irreducible ${\mathbb{F}}_{2}G^{(\infty)}$-module. This deduces that $M\cong S_{{\mathbb{F}}_{2}}$ as ${\mathbb{F}}_{2}T$-modules. Note that $\dim T_{{\mathbb{F}}_{2}}=f\cdot(\binom{n/f}{2}-1)$. By some elementary calculations, we obtain that $\dim T_{{\mathbb{F}}_{2}}=\dim M=n$ has no no integer solutions.
@@
    [9, Lemma 3] $K$ is an elementary abelian $p$-group of order $|{\mathbf{Z}}(N)|^{n}$;
@@
By Lemma 3.1 (i), we have $|K|=|C|^{n}$, and thus $|K|=|\{(c_{1},\dots,c_{n})\mid c_{i}\in C\}|$. Hence elements of $K$ and the tuples $(c_{1},\dots,c_{n})$ with $c_{i}\in C$ are one-to-one correspondent. It follows that elements of the same coset $Cx$ are conjugate under $K$. Hence, two elements $x,y\in N\setminus C$ are conjugate under ${\mathrm{Aut}}(N)$ if and only if the cosets $Cx,Cy$ are conjugate under $G=A/K$. Equivalently, the number of $A$-orbits on $N\setminus C$ is equal to the number of $G$-orbits on non-identity elements of $N/C=V$, which equals $o(V)-1$. Thus, the number of $A$-orbits on $N$ is equal to $o(V)+o(M)-1$, as in part (iii). For $x,y,z\in N$, we notice that $[x,y]=[y,x]^{-1}$ and $[x,yz]=[x,y]^{z}[x,z]=[x,y][x,z]$ as $N^{\prime}={\mathbf{Z}}(N)$. Hence the commutator operator induces an alternative linear map:
@@
Then $\Phi(N)=N^{\prime}N^{2}=N^{\prime}={\mathbf{Z}}(N)$, and $N$ is a special $2$-group. Let $V$ and $M$ be the vector space corresponding to $N/{\mathbf{Z}}(N)$ and ${\mathbf{Z}}(N)$ respectively. Let $G={\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$, the group induced by ${\mathrm{Aut}}(N)$ on the factor group $N/\Phi(N)$. Lemma 3.1 shows that $V$ and $M$ are ${\mathbb{F}}_{2}G$-modules for $G$, and $M$ is ${\mathbb{F}}_{2}G$-isomorphic to a quotient of $\Lambda_{{\mathbb{F}}_{2}}^{2}(V)$. First, we show that $G$ acts transitively on non-zero vectors of both $V$ and $M$. Recall that ${\mathbf{Z}}(N)=N^{\prime}$ consists of all involutions of $N$ and the identity. Hence, non-zero vectors of $M$ are in the same orbits of ${\mathrm{Aut}}(N)$. Since $G^{{\mathbf{Z}}(N)}={\mathrm{Aut}}(N)^{{\mathbf{Z}}(N)}$ by Lemma 3.1(ii), it follows that $G$ is transitive on non-zero vectors of $M$. Note that every element in $N\setminus{\mathbf{Z}}(N)$ is of order $4$. Suppose that $x,y\in N\setminus{\mathbf{Z}}(N)$. Then $x^{2}$ and $y^{2}$ are involutions, and hence there exists $\alpha\in{\mathrm{Aut}}(N)$ such that $(x^{2})^{\alpha}=y^{2}$. This yields that $(x^{\alpha})^{2}=y^{2}$. We may decompose $N=x_{1}{\mathbf{Z}}(N)\cup x_{2}{\mathbf{Z}}(N)\cup\cdots\cup x_{|{\mathbf{Z}}(N)|}{\mathbf{Z}}(N)$ for some $x_{i}\in N\setminus{\mathbf{Z}}(N)$ as $|N/{\mathbf{Z}}(N)|=|{\mathbf{Z}}(N)|$. Then, for each representative $x_{i}$ and any element $a\in M$, we have $(ax_{i})^{2}=a^{2}x_{i}^{2}=x_{i}^{2}$. Thus, we have
@@
Suppose that ${\mathrm{Aut}}(N)$ is non-solvable. Then $G$ is non-solvable by Lemma 3.1(i). Since $\dim M=\dim V$, we have that $G^{(\infty)}\cong{\mathrm{SL}}_{3}(2^{f})$ and $M$ is ${\mathbb{F}}_{2}G^{(\infty)}$-isomorphic to the dual $V$ by Theorem 1.2. Let $v=x{\mathbf{Z}}(N)$ be a non-zero vector of $V$ where $x\in N\setminus{\mathbf{Z}}(N)$. Then $x$ is of order $4$, and $w:=x^{2}$ is a non-zero vector in $M$. Note that an element of $G^{(\infty)}$ stabilizing $v$ must stabilizes $w$ since $(ax)^{2}=x^{2}$ for any $a\in{\mathbf{Z}}(N)$, and hence $G^{(\infty)}_{v}\leqslant G^{(\infty)}_{w}$, which contradicts the fact that $M$ is the dual ${\mathbb{F}}_{2}G^{(\infty)}$-module of $V$ for $G^{(\infty)}\cong{\mathrm{SL}}_{3}(2^{f})$. Thus, ${\mathrm{Aut}}(N)$ is solvable, and $N$ is a Suzuki $2$-group by Shaw’s Theorem [21]. The solvability of ${\mathrm{Aut}}(N)$ is proved by Bryukhanova in [3, Theorem 1.2]. ∎

### v2

@@ Abstract.
Building upon previous results, a classification is given of finite $p$-groups whose elements of order $p$ are all fused. This particularly confirms a conjecture of Gross proposed in 1976 on $2$-automorphic $2$-groups, which are $2$-groups with involutions forming a single fusion class. As a consequence, two open problems regarding AT-groups and FIF-groups are solved.
@@ 1. Introduction
It is well-known that finite $p$-groups with a single subgroup of order $p$ are cyclic or a generalized quaternion group [11, page 199]. The class of finite $p$-groups with a single fusion class of elements of order $p$ has been extensively studied. (Recall that a fusion class of a group $N$ is an orbit of ${\mathrm{Aut}}(N)$ on the elements.) The study was initiated by G. Higman [13] in 1963 when he classified finite non-abelian $2$-groups (with more than one involution) that have cyclic groups of automorphisms transitive on the involutions, called Suzuki $2$-groups.
For odd primes $p$, Shult [24] showed that homocyclic $p$-groups are the only $p$-groups of which elements of order $p$ form a single fusion class. Shaw [23] showed that a finite non-abelian $2$-group (with more than one involution) which has a solvable group of automorphisms transitive on the involutions is a Suzuki $2$-group. Higman’s classification associated with Shaw’s work forms Sections 5-7 of Chapter VII of the book written by Huppert and Blackburn [14].
@@
We use $G^{(\infty)}$ for the unique perfect group appearing in the derived series of $G$. The proof of Theorem 1.1 for $2$-groups depends on certain representations of finite transitive linear groups over ${\mathbb{F}}_{2}$, stated below, which is independently interesting.
@@
Solving Gross’ conjecture is also an important part for the study of some other classes of groups. A group $N$ is called an AT-group if ${\mathrm{Aut}}(N)$ is transitive on the elements of the same order. Zhang [27] presents a description of finite AT-groups. However, the problem of determining AT-$2$-groups has been left open, as commented that “If Gross’ conjecture on $2$-automorphic $2$-groups is true, …… We do not know if Suzuki $2$-groups are AT-groups” before Lemma 2.2 of [27]. A group is called an FIF-group if any two elements of the same order are fused or inverse-fused. Praeger and the first-named author [18] give a description of finite FIF-groups, with the problem of classifying FIF-$2$-groups left open, see [18, Problem 1.5]. As a corollary of Theorem 1.1, these problems are solved.
@@
The three families of groups in Corollary 1.3(iv) are exactly all $2$-groups with $3$ fusion classes, which were also obtained by Bors and Glasby in [2]. We remark that the automorphism groups of the groups $A_{2}(n,\theta)$, $B_{2}(n)$ and $P(\epsilon)$ are determined in Lemmas 4.2, 4.4 and 4.7, respectively.
The layout of the paper is as follows. In Section 2, we introduce some methods in representation theory and prove Theorem 1.2. We prove Theorem 1.1 in Section 3 by linking special $p$-groups with representation theory and using Theorem 1.2. The final section contains the proof of Corollary 1.3 together with the constructions of groups in the Corollary.
@@ Acknowledgments
The authors acknowledge the support of NNSFC grant no. 11931005. Moreover, they would like to thank the anonymous referee for their careful reading of the paper and valuable suggestions.
@@ 2. Transitive linear groups and their exterior squares
All fields and groups are finite in this section. For a field ${\mathrm{F}}$ and a group $G$, an ${\mathrm{F}}G$-reprensentation is a group homomorphism $\rho$ from $G$ to ${\mathrm{GL}}(V)$ for some ${\mathrm{F}}$-space $V$. Every ${\mathrm{F}}G$-reprensentation $\rho$ induces an ${\mathrm{F}}G$-module structure on $V$ by $(xg)\cdot v=x\cdot v^{\rho(g)}$ for $x\in{\mathrm{F}}$ and $g\in G$. If ${\mathrm{E}}$ is an extension field of ${\mathrm{F}}$, then $V^{\mathrm{E}}:=V\otimes{\mathrm{E}}$ is naturally an ${\mathrm{E}}G$-module. Some elementary properties of the functor $-\otimes{\mathrm{E}}$ can be found in [14, Section VII.1], [1, Chapter 9] and [15, Chapter 9]. For an ${\mathrm{F}}G$-module $V$ with representation $\rho$, we use the following two notations:
@@
    [14, Chapter VII, 1.16(a)] $(V_{\mathrm{K}})^{\mathrm{F}}\cong\sum_{\sigma\in{\mathrm{Gal}}({\mathrm{F}}/{\mathrm{K}})}V^{\sigma}$;
@@
    [Deuring-Noether, see [14, Chapter VII, 1.22]] $V^{\mathrm{E}}\cong W^{\mathrm{E}}$ if and only if $V\cong W$.
@@
Let $V={\mathbb{F}}_{2}^{n}$ be a vector space, and let $G\leqslant{\mathrm{GL}}(V)\cong{\mathrm{GL}}_{n}(2)$ be transitive on $V\setminus\{0\}$. These groups $G$ are known by the classification of the finite $2$-transitive permutation groups, and we list them in the next theorem, see [5, Table 7.3].
@@
The following lemma can be directly deduced by Steinberg’s Twisted Tensor Product Theorem [25] (also see [16, Theorem 5.4.1]).
@@
Lemma 2.7 ([19, Theorem 2.2]).
Let $U={\mathbb{F}}_{q}^{m}$, and let $G={\mathrm{SL}}_{m}(q)$. Then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is an irreducible ${\mathbb{F}}_{q}G$-module which cannot be written over any subfield of ${\mathbb{F}}_{q}$.
@@
Results of exterior square of the natural modules of symplectic groups can be found in [6, 17, 21], and the result of the exceptional group of Lie type ${\mathrm{G}}_{2}(q)<{\mathrm{Sp}}_{6}(q)$ (even $q$) can be deduced from [22, Theorem 8.14].
@@
Suppose that $U={\mathbb{F}}_{q}^{2m}$ is a symplectic space with $q=2^{f}$, and $G={\mathrm{Sp}}_{2m}(q)$ or ${\mathrm{G}}_{2}(q)$($m=3$). Then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ has a maximal ${\mathbb{F}}_{q}G$-submodule $T$ of codimension $1$. Moreover, $T$ is indecomposable and one the following statements holds:
@@
Let $G\leqslant{\mathrm{GL}}(V)$ with $V={\mathbb{F}}_{2}^{n}$, and let $M=\Lambda^{2}_{{\mathbb{F}}_{2}}(V)/W$ for a submodule $W$. Assume that $G$ is transitive on $V\setminus\{0\}$. Then $G$ is listed in Theorem 2.4. If $G$ is one of the small groups listed in Theorem 2.4(ii), then computations in Magma [3] shows that $\dim M=1$, which is a contradiction as $\dim V=\dim M=1$ implies $G={\mathrm{GL}}_{1}(2)=1$. Thus, we only need to focus on the cases that $G^{(\infty)}\cong{\mathrm{SL}}_{m}(q)$ ($m\geqslant 2$ and $(m,f)\neq(2,1)$), ${\mathrm{Sp}}_{2m}(q)$ ($m\geqslant 2$ and $(m,q)\neq(2,2)$) or ${\mathrm{G}}_{2}(q)$ ($q\geqslant 4$) for $q=2^{f}$.
Let $K$ be the kernel of $G$ acting on $M$. Then $G/K$ is faithful and transitive on $M\setminus\{0\}$, and either $G^{(\infty)}\leqslant K$ or $K\cap G^{(\infty)}\leqslant{\mathbf{Z}}(G^{(\infty)})$. Suppose that $G^{(\infty)}\leqslant K$. Then $G/K$ is isomorphic to a quotient group of $G/G^{(\infty)}\lesssim{\Gamma\mathrm{L}}(1,2^{f})$. So
@@
We thus have that $K\cap G^{(\infty)}\leqslant{\mathbf{Z}}(G^{(\infty)})$. So $G/K$ is faithful and transitive on $M\setminus\{0\}$, and has a simple composition factor isomorphic to one of ${\mathrm{PSL}}_{m}(q)$ ($m\geqslant 2$ and $(m,f)\neq(2,1)$), ${\mathrm{PSp}}_{2m}(q)$ ($m\geqslant 2$ and $(m,q)\neq(2,2)$) and ${\mathrm{G}}_{2}(q)$ ($q\geqslant 4$) for even $q$. Note that none of these groups is isomorphic to ${\mathrm{A}}_{6}$, ${\mathrm{A}}_{7}$ or ${\mathrm{PSU}}_{3}(3)$, inspecting the candidates for $G/K$ in Theorem 2.4, we conclude that $(G/K)^{(\infty)}\cong G^{(\infty)}$. It follows that $K\cap G^{(\infty)}=1$ and $G^{(\infty)}$ is transitive on $M\setminus\{0\}$, and further $G^{(\infty)}$ is also transitive on $V\setminus\{0\}$.
@@
Lemma 2.6 deduces that each $B_{j}$ is an irreducible ${\mathbb{F}}_{2}G^{(\infty)}$-module. Suppose that $B_{j}$ is not a subspace of $W$ for some $j$. Then $M\cong B_{j}$ as ${\mathbb{F}}_{2}G^{(\infty)}$-modules. It follows that $\dim M=\dim B_{j}$. Note that $\dim B_{j}=n^{2}/f$ for $j\neq f/2$ and $\dim B_{f/2}=n^{2}/(2f)$ when $f$ is even. Hence, $M\cong B_{j}$ only when $j=f/2$, $f=n/2$ and $G^{(\infty)}\cong{\mathrm{SL}}_{2}(2^{f})$. In this case, the induced linear group of $G^{(\infty)}$ on $B_{f/2}$ is isomorphic to $\Omega^{-}_{4}(2^{f/2})$ (see [16, page 45]), which cannot be transitive on non-zero vectors of ${\mathbb{F}}_{2}^{2f}$. Thus, $B_{j}\leqslant W$ for each $j$, and then $M$ is ${\mathbb{F}}_{2}G^{(\infty)}$ isomorphic to a quotient of $A\cong(\Lambda^{2}_{{\mathbb{F}}_{q}}(U))_{{\mathbb{F}}_{2}}$.
Suppose that $G^{(\infty)}\cong{\mathrm{Sp}}_{2m}(q)$ for $m\geqslant 2$ (or $G^{(\infty)}\cong G_{2}(q)$) with $q=2^{f}$. Let $T$ and $T_{0}$ be ${\mathbb{F}}_{q}G^{(\infty)}$-submodules of $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ give in Lemma 2.8. Assume that $m$ is even. Then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is indecomposable with $T$ a maximal ${\mathbb{F}}_{q}G^{(\infty)}$-submodule by Lemma 2.8. Note that $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)/T$ is a trivial ${\mathbb{F}}_{q}G^{(\infty)}$-module. This yields that $M$ is a trivial ${\mathbb{F}}_{2}G^{(\infty)}$-module, a contradiction. Assume that $m$ is odd. Then $T$ is irreducible and cannot be written over any subfield by Lemma 2.8. Hence, $T_{{\mathbb{F}}_{2}}$ is an irreducible ${\mathbb{F}}_{2}G^{(\infty)}$-module. This deduces that $M\cong S_{{\mathbb{F}}_{2}}$ as ${\mathbb{F}}_{2}T$-modules. Note that $\dim T_{{\mathbb{F}}_{2}}=f\cdot(\binom{n/f}{2}-1)$. By some elementary calculations, we obtain that $\dim T_{{\mathbb{F}}_{2}}=\dim M=n$ has no no integer solutions.
@@
    [10, Lemma 3] $K$ is an elementary abelian $p$-group of order $|{\mathbf{Z}}(N)|^{n}$;
@@
By Lemma 3.1 (i), we have $|K|=|C|^{n}$, and thus $|K|=|\{(c_{1},\dots,c_{n})\mid c_{i}\in C\}|$. Hence elements of $K$ and the tuples $(c_{1},\dots,c_{n})$ with $c_{i}\in C$ are one-to-one correspondent. It follows that elements of the same coset $Cx$ are conjugate under $K$. Hence, two elements $x,y\in N\setminus C$ are conjugate under ${\mathrm{Aut}}(N)$ if and only if the cosets $Cx,Cy$ are conjugate under $G=A/K$. Equivalently, the number of $A$-orbits on $N\setminus C$ is equal to the number of $G$-orbits on non-identity elements of $N/C=V$, which equals $o(V)-1$. Thus, the number of $A$-orbits on $N$ is equal to $o(V)+o(M)-1$, as in part (iii).
For $x,y,z\in N$, we notice that $[x,y]=[y,x]^{-1}$ and $[x,yz]=[x,y]^{z}[x,z]=[x,y][x,z]$ as $N^{\prime}={\mathbf{Z}}(N)$. Hence the commutator operator induces an alternating linear map:
@@
Then $\Phi(N)=N^{\prime}N^{2}=N^{\prime}={\mathbf{Z}}(N)$, and $N$ is a special $2$-group. Let $V$ and $M$ be the vector space corresponding to $N/{\mathbf{Z}}(N)$ and ${\mathbf{Z}}(N)$ respectively. Let $G={\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$, the group induced by ${\mathrm{Aut}}(N)$ on the factor group $N/\Phi(N)$. Lemma 3.1 shows that $V$ and $M$ are ${\mathbb{F}}_{2}G$-modules for $G$, and $M$ is ${\mathbb{F}}_{2}G$-isomorphic to a quotient of $\Lambda_{{\mathbb{F}}_{2}}^{2}(V)$.
First, we show that $G$ acts transitively on non-zero vectors of both $V$ and $M$. Recall that ${\mathbf{Z}}(N)=N^{\prime}$ consists of all involutions of $N$ and the identity. Hence, non-zero vectors of $M$ are in the same orbits of ${\mathrm{Aut}}(N)$. Since $G^{{\mathbf{Z}}(N)}={\mathrm{Aut}}(N)^{{\mathbf{Z}}(N)}$ by Lemma 3.1(ii), it follows that $G$ is transitive on non-zero vectors of $M$. Note that every element in $N\setminus{\mathbf{Z}}(N)$ is of order $4$. Suppose that $x,y\in N\setminus{\mathbf{Z}}(N)$. Then $x^{2}$ and $y^{2}$ are involutions, and hence there exists $\alpha\in{\mathrm{Aut}}(N)$ such that $(x^{2})^{\alpha}=y^{2}$. This yields that $(x^{\alpha})^{2}=y^{2}$. We may decompose $N=x_{1}{\mathbf{Z}}(N)\cup x_{2}{\mathbf{Z}}(N)\cup\cdots\cup x_{|{\mathbf{Z}}(N)|}{\mathbf{Z}}(N)$ for some $x_{i}\in N\setminus{\mathbf{Z}}(N)$ as $|N/{\mathbf{Z}}(N)|=|{\mathbf{Z}}(N)|$. Then, for each representative $x_{i}$ and any element $a\in M$, we have $(ax_{i})^{2}=a^{2}x_{i}^{2}=x_{i}^{2}$. Thus, we have
@@
Suppose that ${\mathrm{Aut}}(N)$ is non-solvable. Then $G$ is non-solvable by Lemma 3.1(i). Since $\dim M=\dim V$, we have that $G^{(\infty)}\cong{\mathrm{SL}}_{3}(2^{f})$ and $M$ is ${\mathbb{F}}_{2}G^{(\infty)}$-isomorphic to the dual $V$ by Theorem 1.2. Let $v=x{\mathbf{Z}}(N)$ be a non-zero vector of $V$ where $x\in N\setminus{\mathbf{Z}}(N)$. Then $x$ is of order $4$, and $w:=x^{2}$ is a non-zero vector in $M$. Note that an element of $G^{(\infty)}$ stabilizing $v$ must stabilizes $w$ since $(ax)^{2}=x^{2}$ for any $a\in{\mathbf{Z}}(N)$, and hence $G^{(\infty)}_{v}\leqslant G^{(\infty)}_{w}$, which contradicts the fact that $M$ is the dual ${\mathbb{F}}_{2}G^{(\infty)}$-module of $V$ for $G^{(\infty)}\cong{\mathrm{SL}}_{3}(2^{f})$. Thus, ${\mathrm{Aut}}(N)$ is solvable, and $N$ is a Suzuki $2$-group by Shaw’s Theorem [23]. The solvability of ${\mathrm{Aut}}(N)$ is proved by Bryukhanova in [4, Theorem 1.2]. ∎